# Conformal Prediction Assessment: A Framework for Conditional Coverage Evaluation and Selection

### 28 Mar 2026 v1

@@ 4.2.1 Ranking Consistency Analysis
Table 2 demonstrates the framework’s robust discriminative power across diverse distributional regimes. In heteroscedastic environments characterized by significant variation in local coverage (Setting C), CPA exhibits high concordance with the Oracle ranking ($\tau_{w}>0.9$). Crucially, in high-noise or highly correlated feature scenarios (Settings B & D), we observe a divergence between exact ordering and selection utility: while stochastic perturbations may permute the ranking of statistically indistinguishable mid-tier models (yielding moderate $\tau_{w}$), the consistently high NDCG scores demonstrate that CPA successfully identifies the optimal candidate. Even in the context of performance saturation (Setting A)—where asymptotic equivalence among methods reduces the Hit@3 metric—the high global correlation confirms that the framework correctly separates the cluster of high-performing models from suboptimal alternatives.
@@ 4.2.2 Distributional Recovery via CVP
Beyond scalar metrics, we examine the distributional convergence of the estimated reliability by comparing the Oracle CVP—the empirical quantile function of true coverage probabilities—against the CPA estimates. As illustrated in Figure 4 (Setting B), the estimated CVP tracks the Oracle distribution with high fidelity. The framework successfully reconstructs the distinctive reliability profile characteristic of rigid methods (e.g., CP-Residual)—indicative of over-coverage on easy instances versus collapse on outliers—while distinctively mapping the calibrated profiles of adaptive strategies (e.g., CQR and CP-Studentized). This confirms that CPA captures the full distributional characteristics of the uncertainty distribution, offering a more granular diagnosis than aggregate coverage statistics.
@@ 4.3 Robustness to Reliability Estimator Misspecification
A critical consideration for practical deployment is the sensitivity of assessment outcomes to the specification of the reliability estimator $\hat{\eta}$. To systematically investigate this, we benchmarked our recommended AutoML Baseline (an ensemble with Isotonic Calibration) against six perturbations designed to probe specific failure modes. These range from uncalibrated variants to models subject to severe structural misspecification (see Appendix G for detailed specifications).
Figure 5 visualizes the impact of these perturbations on ranking fidelity. The results reveal a distinct bifurcation in performance metrics:
  • Stability of Selection Utility: While CPA consistently identifies top models (high NDCG@1), the degradation of Logistic Regression in nonlinear regimes (e.g., Setting B) highlights the necessity of sufficient model capacity for complex boundaries.
  • Monotonicity Drives Selection: The similar NDCG in most settings between the Baseline and Uncalibrated configurations implies that monotonicity is the primary determinant of selection utility. While strict probabilistic calibration remains requisite for accurate absolute auditing, it is secondary for the relative ordering of candidates.
However, a distinct failure mode is identified in the nonlinear Setting B, where the Logistic Regression estimator exhibits a severe breakdown in rank correlation. As detailed in the bias analysis in Appendix H, this failure is attributable to structural misspecification: the linear decision boundary lacks the flexibility to model the complex level sets of the true conditional coverage probability. This highlights a key prerequisite: the reliability estimator must have sufficient model complexity.
@@ 4.4 Discussion of Simulation Results
The simulation studies validate CPA as a robust diagnostic instrument for auditing uncertainty quantification, demonstrating the CVI metric and CVP curves as consistent estimators of ground-truth performance. Based on these findings, we recommend employing an AutoML-based ensemble with Isotonic Calibration as the default reliability estimator for robustness to model mismatch. Regarding data allocation, our sensitivity analysis (detailed in Appendix I) indicates that a balanced split ratio ($\rho=0.5$) mitigates the trade-off between approximation and estimation errors. Furthermore, reliable auditing requires sufficient data to learn the failure boundary; our empirical investigation suggests a minimum sample size of $N\approx 800$ for standard regression tasks (see Appendix J).
@@ 5.1.1 Benchmark Datasets
Table 3 summarizes the nine benchmark datasets used in our evaluation, including their scale, dimensionality, prediction target, and application domain(Romano et al., 2019; Agarwal et al., 2025). All tasks are standard real-valued regression problems.
@@
Miami_House $13,932$ 15 Sale price Real Estate
@@ 5.1.2 Experimental Protocol
For each dataset, we adopt a nested data-splitting design to separate evaluation from final performance assessment. In $10$ independent repetitions, the data are divided into an 80% Master Train set and a 20% Master Test set, with the latter held out for final assessment.
Within the Master Train set, we perform $K=5$ random 50/50 splits to train candidate conformal predictors and estimate reliability. In each split $k$, models are trained on an Internal Train subset and evaluated on an Internal Validation subset, producing coverage indicators $I_{i}=\mathbbm{1}\{Y_{i}\in\widehat{\mathcal{C}}(X_{i})\}$. These indicators serve as labels for fitting the reliability estimator $\hat{\eta}^{(k)}(x)$, which is then used by CC-Select for model choice. For final evaluation, the selected predictor is re-trained on the full Master Train set and assessed on the Master Test set. The reliability score for a test point $X_{j}$ is obtained by ensembling the inner estimators: $\hat{\eta}(X_{j})=\frac{1}{K}\sum_{k=1}^{K}\hat{\eta}^{(k)}(X_{j}).$ These scores support the construction of reliability diagrams and conditional performance metrics on the held-out Master Test data.
@@ 5.2 Reliability Checking
We assess the calibration of the reliability estimator $\hat{\eta}(x)$ to ensure that the predicted probabilities accurately reflect true coverage frequencies. We employ two standard diagnostics: Reliability Diagrams, which visualize the alignment between predicted probability and empirical coverage relative to the identity line $y=x$, and the Expected Calibration Error (ECE), which quantifies the weighted average absolute deviation from perfect calibration. The formal definitions and detailed methodology are provided in Appendix C.
Results.   Figure 6 shows that $\hat{\eta}(x)$ is well calibrated across all nine datasets: empirical coverage closely matches predicted reliability with ECE values typically below 0.01. This supports using CPA to audit conformal predictors and validates $\hat{\eta}(x)$ as a practical proxy for conditional coverage.
Figure 6: Reliability diagrams for CPA across nine datasets. Each plot compares empirical coverage (y-axis) to predicted reliability (x-axis) using $B=10$ equal-frequency bins; the diagonal indicates perfect calibration.
@@ 5.3 Model Selection
We assess the efficacy of the CC-Select algorithm (Algorithm 2) in identifying conformal prediction procedures that optimize conditional coverage.
@@ 5.3.1 Experimental Setup and Metrics
Performance on the held-out Master Test set is evaluated using three complementary metrics: Marginal Coverage (target $1-\alpha=0.9$), Average Interval Length (predictive efficiency), and the Worst-Slab Coverage (WSC) (Cauchois et al., 2021). WSC, a practical finite-sample proxy for conditional validity, measures the minimum empirical coverage across disjoint strata of the covariate space. Details on WSC are provided in Appendix K.1.
@@ 5.3.2 Results
Figure 7 displays the distribution of WSC scores across the nine datasets. The procedure selected by CC-Select (labeled “Selected”) consistently achieves high WSC scores, often outperforming individual candidates. Detailed numerical results, including mean interval lengths and marginal coverage levels, are available in Table 5 in Appendix K.2. It is important to note that, on certain datasets, some CP methods can perform extremely poorly (e.g., on the Meps_21 dataset). This observation highlights the necessity of CPA-type tools for systematically auditing CP methods, in order to detect and prevent inadequate or even severely miscalibrated coverage behavior. Collectively, these empirical results confirm that the CVI metric serves as an effective, data-driven criterion for auditing and selecting safe conformal predictors in complex real-world environments.
@@
5.4 Diagnosing Heterogeneity in Bike Sharing Demand
This case study illustrates CPA’s capacity as a diagnostic tool to uncover covariate-dependent prediction failure modes, which are obscured by marginal metrics. We focus on the Bike dataset, where the target is the hourly rental count, and covariates include temporal, calendar, and weather variables. We apply the nested protocol (Section 5.1) and examine three representative conformal procedures: CP-Residual, CP-Studentized, and CQR. We also conduct experiments examining interactions between two covariates, as detailed in Appendix K.3.
We examine how conditional coverage $\hat{\eta}(x)$ varies along individual covariates. Figure 8 (top) plots the estimated coverage as a function of the hour of the day. Although all methods achieve near-nominal marginal coverage, heterogeneity is pronounced: the residual-based method exhibits noticeable undercoverage during the commuting peaks (8–9am and 5–6pm), a high-volatility regime where the homoscedastic residual model is misspecified. In contrast, CQR maintains coverage closer to the nominal level during these periods. A similar analysis for temperature (Figure 8, bottom) shows that while coverage is satisfactory at moderate temperatures, residual-based methods tend to over-cover at temperature extremes, while CQR remains closer to the nominal level. These univariate analyses expose systematic coverage deficits in specific regions (e.g., rush hours) invisible from global metrics.
@@ 6 Conclusion
Practically, CPA offers a principled alternative to binning-based diagnostics and demonstrates how conditional validity can guide model selection through the CC-Select procedure. The learned reliability scores may also be used as deployment-time indicators for identifying regions of unreliable uncertainty quantification.
@@
Figure 8: Univariate reliability patterns on the Bike dataset. Top: Conditional coverage vs. hour. Bottom: Conditional coverage vs. temperature. Blue bars show empirical coverage; orange lines show estimated reliability $\hat{\eta}(X)$ with standard deviation bands.
@@
  • Algorithmic Specifications (Appendix B). Provides the formal pseudocode for the core components of our framework, specifically the Ensemble Reliability Estimator Training (CPA-Train) and the selection workflow (CC-Select).
  • Experimental Setup (Appendices DE). Delineates the reproducibility details for the simulation studies. Appendix D specifies the mathematical formulas for the four synthetic Data Generating Processes (DGPs). Appendix E details the implementation, hyperparameters, and base learners for all nine benchmarked uncertainty quantification methods.
  • Metric Definitions (Appendix F). Formally defines the ranking metrics used to evaluate selection fidelity, including Weighted Kendall’s $\tau_{w}$, NDCG, and Hit@k.
  • Robustness and Failure Mechanisms (Appendices GH). Supports the robustness analysis in Section 4.3. Appendix G lists the exact configurations for the reliability estimator perturbations. Appendix H provides a deep-dive investigation into the mechanism of failure (Bias Analysis), elucidating why linear estimators fail in nonlinear regimes.
  • Extended Real-Data Analysis (Appendix K). Expands upon the real-world experiments. It includes the formal definition of the Worst Slab Coverage (WSC) metric and provides the full table of numerical results (average length, marginal coverage, and WSC) for all datasets.
@@ Appendix A Proofs and Illustrations of the Main Theorems
A.1 Proof of Theorem 1
@@ A.3 Proof of Theorem 3
The desired conclusion thus follows. ∎
@@ A.4 Proof of Theorem 4
Appendix B Detailed Algorithms

### 23 Apr 2026 v2

@@ 4.2.1 Ranking Consistency Analysis
Table 2 demonstrates that CPA retains strong discriminative power across diverse regimes. In the heteroscedastic setting, where local coverage varies substantially (Setting C), CPA closely matches the Oracle ranking ($\tau_{w}>0.9$). In the high-noise and strongly correlated settings (Settings B and D), exact ordering and selection utility diverge: stochastic fluctuations can permute statistically similar mid-ranked methods, which lowers $\tau_{w}$, but the uniformly high NDCG scores show that CPA still identifies the best candidate reliably. Even in the saturation regime (Setting A), where several methods perform nearly identically and Hit@3 is therefore less informative, the high global correlation indicates that CPA correctly separates the strong-performing methods from clearly inferior ones.
@@ 4.2.2 Distributional Recovery via CVP
Beyond scalar metrics, we also examine distributional recovery by comparing the estimated reliability profile with the Oracle CVP, the empirical quantile function of the true conditional coverage probabilities. Figure 3 for Setting B shows that the estimated CVP closely tracks the Oracle distribution. CPA recovers the characteristic profile of rigid methods such as CP-Residual, which tend to over-cover easy points while failing on outliers, and it also preserves the more balanced profiles of adaptive methods such as CQR and CP-Studentized. These results indicate that CPA captures the full distribution of reliability rather than only an aggregate summary, enabling a more granular diagnosis than marginal coverage alone.
@@ 4.3 Robustness to Reliability Estimator Misspecification
A critical practical question is how sensitive CPA is to the specification of the reliability estimator $\hat{\eta}$. To study this systematically, we compare our recommended AutoML Baseline—an ensemble with isotonic calibration—against six perturbations chosen to expose specific failure modes. These variants range from uncalibrated estimators to models with severe structural misspecification; Appendix H provides the full specifications.
Figure 4 visualizes how these perturbations affect ranking fidelity. Two patterns emerge:
  • Selection utility is stable. CPA usually identifies the top-performing models, as reflected by the consistently high NDCG@1 values. The main exception is Logistic Regression in nonlinear regimes such as Setting B, where insufficient model capacity limits performance.
  • Monotonicity primarily drives ranking. In most settings, the Baseline and Uncalibrated estimators achieve similar NDCG values, suggesting that preserving the correct ordering of reliability is more important for model selection than perfect probabilistic calibration. Calibration remains essential for accurate absolute auditing, but it is secondary for relative ranking.
However, nonlinear Setting B exposes a clear failure mode: the Logistic Regression estimator suffers a severe drop in rank correlation. As the bias analysis in Appendix I explains, this breakdown is caused by structural misspecification: a linear decision boundary cannot capture the complex level sets of the true conditional coverage function. This finding highlights a key practical prerequisite for CPA: the reliability estimator must have sufficient model complexity.
@@ 4.4 Discussion of Simulation Results
Overall, the simulation results validate CPA as a robust tool for auditing uncertainty quantification. The CVI metric and CVP curves consistently recover the ground-truth reliability landscape, both for ranking and for distributional diagnosis. Based on these findings, we recommend an AutoML-based ensemble with Isotonic Calibration as the default reliability estimator because it is robust to model mismatch. For data allocation, the sensitivity analysis in Appendix J suggests that a balanced split ratio ($\rho=0.5$) best balances approximation and estimation error. Finally, reliable auditing requires enough data to learn the failure boundary; our empirical results suggest a minimum sample size of roughly $N\approx 800$ for standard regression tasks (Appendix K).
@@ 5.1.1 Benchmark Datasets
Table 3 summarizes the nine benchmark datasets used in our evaluation, which are drawn from the real-data benchmarks considered in Romano et al. (2019) and Agarwal et al. (2025). This choice aligns our empirical study with established evaluation protocols in the literature, rather than relying on arbitrarily selected datasets. The table reports their scale, dimensionality, prediction target, and application domain. All tasks are standard real-valued regression problems.
@@
Miami_2016 $13,932$ 15 Sale price Real Estate
@@ 5.1.2 Experimental Protocol
For each dataset, we use a nested data-splitting design to separate model assessment from final evaluation. Across $10$ independent repetitions, the data are split into an 80% Master Train set and a 20% Master Test set, with the latter reserved for final assessment.
Within the Master Train set, we perform $K=5$ random 50/50 splits to train candidate conformal predictors and estimate reliability. In split $k$, models are fitted on an Internal Train subset and evaluated on an Internal Validation subset, yielding coverage labels $I_{i}=\mathbbm{1}\{Y_{i}\in\widehat{\mathcal{C}}(X_{i})\}$ for fitting the reliability estimator $\hat{\eta}^{(k)}(x)$. CC-Select uses these estimators for model choice. For final evaluation, the selected predictor is retrained on the full Master Train set and assessed on the Master Test set, while the reliability score for a test point $X_{j}$ is computed by ensembling the inner estimators, $\hat{\eta}(X_{j})=\frac{1}{K}\sum_{k=1}^{K}\hat{\eta}^{(k)}(X_{j})$. These scores are used to construct reliability diagrams and conditional performance metrics on the held-out test data.
@@ 5.2 Reliability Checking
We assess the calibration of the reliability estimator $\hat{\eta}(x)$ using two standard diagnostics: Reliability Diagrams, which compare predicted probabilities with empirical coverage, and the Expected Calibration Error (ECE), which measures the weighted average absolute deviation from perfect calibration. Formal definitions and implementation details are provided in Appendix D.
Results.   Figure 5 shows that $\hat{\eta}(x)$ is well calibrated across all nine datasets: empirical coverage closely tracks predicted reliability, with ECE values typically below 0.01. Although these diagnostics do not establish learnability in full generality, they support that the learned reliability surface is a reliable practical proxy for conditional coverage, validating the use of CPA for auditing conformal predictors.
Figure 5: Reliability diagrams for CPA across nine datasets. Each plot compares empirical coverage (y-axis) with predicted reliability (x-axis) using $B=10$ equal-frequency bins.
@@ 5.3 Model Selection
We assess the efficacy of the CC-Select algorithm (Algorithm 2) in identifying conformal prediction procedures that optimize conditional coverage. This is the operational step through which CPA turns diagnostics into action when multiple plausible refinements are available.
@@ 5.3.1 Experimental Setup and Metrics
Performance on the held-out Master Test set is evaluated using three complementary metrics: Marginal Coverage (target $1-\alpha=0.9$), Average Interval Length (predictive efficiency), and the Worst-Slab Coverage (WSC) (Cauchois et al., 2021). WSC, a practical finite-sample proxy for conditional validity, measures the minimum empirical coverage across disjoint strata of the covariate space. Details on WSC are provided in Appendix L.1.
@@ 5.3.2 Results
Figure 6 displays the distribution of WSC scores across the nine datasets. The procedure selected by CC-Select (labeled “Selected”) consistently achieves high WSC scores, often outperforming individual candidates. Detailed numerical results, including mean interval lengths and marginal coverage levels, are available in Table 6 in Appendix L.2. It is important to note that, on certain datasets, some CP methods can perform extremely poorly (e.g., on the Meps_21 dataset). This observation highlights the necessity of CPA-type tools for systematically auditing CP methods, in order to detect and prevent inadequate or even severely miscalibrated coverage behavior. Collectively, these empirical results confirm that the CVI metric serves as an effective, data-driven criterion for auditing and selecting safe conformal predictors in complex real-world environments.
5.4 Practical Utility of CPA: Diagnosis, Refinement, and Selection
CPA is useful not only for auditing conditional validity, but also for diagnosing failure modes of a conformal predictor and guiding the choice among practical refinements. The framework identifies where an existing method fails and compares alternative conformal pipelines through their estimated reliability patterns.
We illustrate this role on the Bike dataset, where the target is hourly rental count and the covariates include temporal, calendar, and weather variables. We take CP-Residual as the baseline conformal predictor and consider two more adaptive variants, CP-Studentized and CQR. The goal is to show how CPA diagnoses the baseline method and helps select a more conditionally valid alternative.
We begin with CP-Residual. Although it attains marginal coverage close to the nominal level, its Conditional Validity Profile in Figure 7(a) reveals substantial heterogeneity in conditional coverage, with both undercoverage and overcoverage across the population. The issue is therefore not marginal calibration, but poor conditional validity across the feature space. To localize this heterogeneity, we examine reliability as a function of hour. Figure 7(b) shows clear undercoverage during peak commuting hours, where a homoscedastic residual score is inadequate.
(a) CVP curve for CP-Residual.
(b) Conditional coverage for CP-Residual.
Figure 7: Diagnostics for the baseline conformal predictor CP-Residual on the Bike dataset. Left: the Conditional Validity Profile (CVP) reveals substantial heterogeneity in conditional coverage despite near-nominal marginal coverage. Right: the hourly reliability pattern shows clear undercoverage during peak commuting hours, and the estimated curve broadly follows the empirical trend.
This diagnosis naturally motivates more adaptive score constructions. A failure pattern concentrated in high-variability regimes suggests methods that better adapt interval width to local uncertainty. CP-Studentized rescales residuals by a local variability estimate, whereas CQR constructs intervals through conditional quantile estimation. Figure 8 shows that both refinements yield CVP curves closer to the nominal target than CP-Residual, indicating improved conditional validity at the population level. The same pattern appears at the feature level: in Figure 9, both methods reduce the undercoverage during peak commuting hours, with CQR exhibiting the most stable hourly reliability pattern.
(a) CP-Studentized
(b) CQR
Figure 8: Conditional Validity Profile (CVP) curves for two adaptive refinements on the Bike dataset. Compared with CP-Residual, both CP-Studentized and CQR exhibit improved conditional validity, with CVP curves closer to the nominal target.
Accordingly, CPA ranks the three methods as CQR first, CP-Studentized second, and CP-Residual last on this dataset. This example makes the practical role of CPA concrete: it reveals where a baseline conformal predictor fails, motivates targeted refinements, and supports the selection of a more reliable conformal method. Additional two-dimensional diagnostics over temp and hour are reported in Appendix L.3.
(a) CP-Studentized
(b) CQR
Figure 9: Conditional coverage by hour for two adaptive refinements on the Bike dataset. Compared with CP-Residual, both CP-Studentized and CQR reduce the undercoverage during peak commuting hours, with CQR exhibiting the most stable hourly pattern.
@@ 6 Conclusion
Practically, CPA offers a principled alternative to binning-based diagnostics and demonstrates how conditional validity can guide model selection through the CC-Select procedure. The learned reliability scores may also be used as deployment-time indicators for identifying regions of unreliable uncertainty quantification. More broadly, CPA is most useful when it helps practitioners improve existing conformal pipelines through diagnosis-guided refinement.
@@
  • Feasibility of Learning the Conditional Coverage Function (Appendix B). Explains why the conditional coverage surface induced by a fixed conformal predictor can be substantially simpler than the full conditional law in practically relevant regimes, and provides supporting synthetic and real-data evidence.
  • Algorithmic Specifications (Appendix C). Provides the formal pseudocode for the core components of our framework, specifically the Ensemble Reliability Estimator Training (CPA-Train) and the selection workflow (CC-Select).
  • Experimental Setup (Appendices EF). Delineates the reproducibility details for the simulation studies. Appendix E specifies the mathematical formulas for the four synthetic Data Generating Processes (DGPs). Appendix F details the implementation, hyperparameters, and base learners for all nine benchmarked uncertainty quantification methods.
  • Metric Definitions (Appendix G). Formally defines the ranking metrics used to evaluate selection fidelity, including Weighted Kendall’s $\tau_{w}$, NDCG, and Hit@k.
  • Robustness and Failure Mechanisms (Appendices HI). Supports the robustness analysis in Section 4.3. Appendix H lists the exact configurations for the reliability estimator perturbations. Appendix I provides a deep-dive investigation into the mechanism of failure (Bias Analysis), elucidating why linear estimators fail in nonlinear regimes.
  • Extended Real-Data Analysis (Appendix L). Expands upon the real-world experiments. It includes the formal definition of the Worst Slab Coverage (WSC) metric and provides the full table of numerical results (average length, marginal coverage, and WSC) for all datasets.
@@ Appendix A Proofs and Illustrations of the Main Theorems
A.1 Illustration of the Limiting Conditional Coverage
This subsection provides illustrative examples to complement Theorem 1. Figure 10 visualizes the asymptotic conditional coverage function $\eta(x)$ under several representative forms of model misspecification, illustrating that uniform convergence of $\eta_{n}(x)$ does not generally imply convergence to the nominal target level $1-\alpha$.
A.2 Proof of Theorem 1
@@ A.4 Proof of Theorem 3
The desired conclusion thus follows.
@@ A.5 Proof of Theorem 4
Appendix B Feasibility of Learning the Conditional Coverage Function
The proposed CPA crucially relies on the estimation of the conditional coverage probability. In full generality, estimating the conditional coverage function of an arbitrary CP method need not be easier than learning the full conditional law of $Y\mid X=x$. For a generic interval-valued predictor $C(x)=[l(x),u(x)]$, one has
$\eta(x)=\mathbb{P}(Y\in C(x)\mid X=x)=F_{Y\mid X}(u(x)\mid x)-F_{Y\mid X}(l(x)\mid x).$
Hence $\eta(x)$ is a functional of the conditional distribution evaluated at the two interval endpoints. For an arbitrary conformal prediction method, estimating $x\mapsto\eta(x)$ may therefore be essentially as difficult as learning the conditional distribution itself.
Fortunately, CPA audits a given conformal predictor rather than attempting to recover the full conditional law of $Y\mid X$. In the residual-score setting analyzed in Section 2.1, Theorem 1 shows that the finite-sample conditional coverage function $\eta_{n}(x)$ converges uniformly to
$\eta(x)=\mathbb{P}\bigl(|\mu(X)+\epsilon-\tilde{\mu}(X)|\leq q_{1-\alpha}\mid X=x\bigr).$
Accordingly, the target of CPA is not an arbitrary functional of $Y\mid X=x$, but the probability of a specific residual-based event induced by the fitted center $\tilde{\mu}(x)$ and the global threshold $q_{1-\alpha}$. This representation makes it possible to see practically relevant regimes in which the induced reliability surface $x\mapsto\eta(x)$ is substantially simpler than the full conditional law itself, as illustrated below.
Example.
Consider the well-specified case $\tilde{\mu}(x)=\mu(x)$. Then
$\eta(x)=\mathbb{P}\bigl(|\epsilon|\leq q_{1-\alpha}\mid X=x\bigr).$
If we denote the conditional distribution function of the residual by
$G_{x}(t):=\mathbb{P}(\epsilon\leq t\mid X=x),$
then
$\eta(x)=G_{x}(q_{1-\alpha})-G_{x}(-q_{1-\alpha}).$
Thus, in the residual-score setting, the target is not the full conditional law of $Y\mid X=x$, but only the probability of a particular residual event. This target can be much simpler whenever the residual law depends on $x$ through a low-dimensional feature, or in the extreme case does not depend on $x$ at all.
A particularly transparent regime is the homoscedastic case $\epsilon\perp X$. Then $G_{x}\equiv G$ does not vary with $x$, so
$\eta(x)=\mathbb{P}(|\epsilon|\leq q_{1-\alpha})$
is constant in $x$. In this regime, a good conformal method induces a trivial reliability surface even when the regression function $\mu(x)$ remains difficult to estimate.
As a concrete example, let $\mu(x)=\|x\|^{\beta}$ for some $0<\beta<1$, let $X$ take values in $1,1]^{d}$, and assume $\epsilon\perp X$. Then $\mu(x)$ is $\beta$-Hölder but non-differentiable at the origin. Classical minimax theory implies that estimating the regression function over a $d$-dimensional Hölder class can be statistically difficult, with a lower bound of order $(n/\log n)^{-\beta/(2\beta+d)}$ under sup-norm loss (Stone, 1982). At the same time, uniformly consistent nonparametric regression estimators still exist under standard conditions (Liero, 1989). The key point for CPA is that the induced target $\eta(x)$ may remain much simpler than the full conditional law even when the regression problem itself is high-dimensional and non-smooth.
To numerically demonstrate this phenomenon, we conducted a high-dimensional simulation in which
$Y=\mu(X)+\sigma(X_{1})\epsilon,\qquad\mu(X)=\|X\|^{1/2},\qquad\sigma(X_{1})=0.5+|X_{1}|,$
with $X\sim\mathrm{Unif}([-1,1]^{200})$ and $\epsilon\sim N(0,1)$ independent of $X$. In this design, the conditional law of $Y\mid X=x$ depends on a high-dimensional, non-smooth regression function, whereas for a fixed residual-score conformal predictor the corresponding conditional coverage function is driven mainly by the one-dimensional feature $X_{1}$. We use $n_{\mathrm{train}}=2000$ observations to construct the conformal predictor and an independent evaluation sample of size $n_{\mathrm{eval}}=500$ to learn $\eta(x)$. On an independent test sample, the pointwise true conditional coverage is approximated by Monte Carlo replication.
Figure 11: High-dimensional illustration of the feasibility claim. Although the ambient dimension is $d=200$, the conditional coverage surface induced by the fixed residual-score conformal predictor varies mainly with $x_{1}$, and a one-dimensional smoother recovers this pattern closely. The MAE and RMSE are computed against Monte Carlo approximations of the pointwise true conditional coverage.
Figure 11 shows that the estimated $\hat{\eta}(x_{1})$ tracks the true conditional coverage curve closely despite the ambient dimension being $d=200$. This experiment is not intended as a universal guarantee. Rather, it provides a concrete illustration that for auditing a fixed conformal predictor, learning the induced reliability surface can be considerably easier than learning the entire conditional law.
We also repeated this experiment for $d\in\{20,50,100,200\}$ while keeping the sample sizes and estimation procedure fixed. Table 4 shows that the MAE and RMSE remain small across dimensions, consistent with the fact that in this example the conditional coverage function is still driven primarily by a low-dimensional feature.
Table 4: Estimation accuracy of the conditional coverage function under different ambient dimensions. Entries report mean (SD) over 10 repetitions.
$d$ 20 50 100 200
MAE 0.026 (0.007) 0.024 (0.004) 0.022 (0.006) 0.026 (0.009)
RMSE 0.034 (0.009) 0.031 (0.005) 0.030 (0.009) 0.033 (0.012)
Finally, the real-data experiments provide further practical support for this perspective. As shown in Figure 5, the learned reliability estimator is well calibrated across the nine benchmark datasets, with empirical coverage closely tracking predicted reliability. Although such real-data evidence cannot deliver the same oracle comparison as the synthetic example above, it does indicate that in realistic applications the learned reliability surface can still serve as a meaningful and empirically well-behaved proxy for conditional coverage assessment.
Appendix C Detailed Algorithms